# FineScope: SAE-guided Data Selection Enables Domain-Specific LLM Pruning & Fine-Tuning

### v2

@@ 3.2.1 Pruning
To enhance model efficiency while retaining domain-specific knowledge, we apply structured pruning using LLM-Pruner (llmpruner, ). Rather than relying on general-purpose criteria, we guide the pruning process with a domain-specific dataset $D_{s}$ (as shown in Figure 2c), so that only components relevant and active to the target domain are preserved. Guided by the domain-specific dataset $D_{s}$, the contribution of each model block is estimated using gradient-based attribution. Specifically, we compute an importance score for each block based on the first-order gradient of the task-specific loss with respect to $D_{s}$. This allows the pruning process to preserve components that are most critical for performance in the target domain.
We apply block-wise structured pruning using LLM-Pruner (llmpruner, ), which removes redundant components of the model and significantly reduces inference cost. The original model, denoted by $\mathcal{M}$, is pruned with respect to the domain-specific dataset $D_{s}$ and a pruning ratio $r$, resulting in a compressed model $\mathcal{M}_{r}$:
@@
3.2.2 Fine-Tuning
Self Distillation Finetuning (SDFT) (sdft, ) is a technique in which a distilled dataset is generated to match the output distribution of the original model, leading to improved generalization. While SDFT is commonly used for knowledge transfer, its role in FineScope extends beyond dataset refinement. We adapt SDFT to address the loss of domain-specific representations caused by pruning. High-confidence predictions from a state-of-the-art teacher model are used to generate a distilled dataset, which is then used to fine-tune the pruned student. This teacher-guided distillation helps restore lost knowledge and improves performance. A high level overview of the process is illustrated in Figure 2d. This approach reduces overfitting to the small curated dataset and compensates for information removed during pruning, even when the teacher and student differ in size or architecture.
In our framework, we adapt SDFT to further refine the pruned model. We generate a distilled dataset using either the original model or a pretrained state-of-the-art teacher, and use it to fine-tune the pruned model. This process helps preserve domain-specific knowledge while improving generalization.
In the original SDFT formulation (sdft, ), given an input $x$, context $c^{t}$, and output $y^{t}$ from the teacher model, the distilled output $y^{\prime}$ is sampled as:
@@ 3.2.2 Fine-Tuning
We adapt this procedure in FineScope by modifying the fine-tuning objective for the pruned model. The resulting self-distillation loss is defined as:
$L_{\text{msdft}}=-\log f_{p}(y^{\prime}\mid c^{t},y^{t}),$ (11)
where $f_{p}$ denotes the pruned model and $L_{\text{msdft}}$ represents the self-distillation loss.
@@ 4 Evaluation
(1) Models. We evaluate our method using three models: Vicuna-7B (vicuna, ), a fine-tuned version of LLaMa 2 (llama2, ); MathCoder-CL-7B (mathcoder, ), a CodeLlama (codellama, ) variant; and LLaMa 3.1-8B (llama3, ).
(2) Baselines. We compare FineScope against six baseline settings: (a) fine tuning the pruned model using randomly selected data of the same size, (b) fine tuning with the full dataset containing mixed domains, (c) fine tuning with Alpaca data using FineScope’s pruning strategy, (d) evaluating pretrained models without any fine tuning, (e) evaluating pretrained models fine tuned with FineScope-curated data, and (f) comparisons against GPT-3 (6.7B and 175B)(gpt3, ) and OLMO-7B(olmo, ) (Table 1), along with GPT-3 (13B and 175B) (Table 2)
@@
(a) Domain Specific Tuning. We prune models using domain specific datasets curated with our SAE-guided framework. The SAEs are trained on the RedPajama dataset (together2023redpajama, ), which includes content from CommonCrawl, C4, GitHub, Wikipedia, Books3, ArXiv, and StackExchange, providing broad domain coverage. Using these SAEs, we curate domain specific subsets from OpenInstruct (haku, ), which aggregates instruction datasets such as Alpaca (albertying, ), Self Instruct (selfinstruct, ), GPT 4 Instruct, Roleplay (GPTeacher, ), Code Alpaca (codealpaca, ), and Dolly (DatabricksBlog2023DollyV2, ). Based on user provided seed samples, we extract curated datasets for STEM (2,100 samples), Social Sciences (2,401 samples), and Humanities (2,374 samples), selecting the most frequently chosen samples across all trained SAEs.
(b) Subdomain Specific Tuning (Math). To assess FineScope’s effectiveness at a more granular level, we evaluate fine tuning on mathematical subdomains. SAEs are trained on the MetaMath dataset (metamath, ), and used to curate subsets from the Math dataset (hendrycksmath2021, ). In this setting, subdomains are merged into a unified pool for curation. From this pool, we extract the Pre Algebra, Algebra, and Counting and Probability subsets. Models are fine tuned using OpenMath2 (openmath, ), and Notus 7B (notus, ) serves as the Alpaca tuned baseline. Evaluations are conducted separately on each subdomain test set (Table 2).
(c) Coding Specific Tuning. Following the same SAE guided curation approach, we construct a code focused dataset from OpenInstruct, resulting in 1,200 examples. To evaluate model performance on code generation, we fine tune on this curated dataset and assess results using the HumanEval (humaneval, ) and MBPP (mbpp, ) benchmarks (Table 3).
(4) Implementation Details SAE training: We ran SAE training (EleutherAI_SAE, ) using the AdamW optimizer with a learning rate of 1e-5, a batch size of $8$, and a Top-K value of $128$. We used GPT-4 (gpt4, ) to generate the $10$ user-defined seeds. Finetuning: We fine-tuned LMs (LORA-fine tuning (lora, )) using the same AdamW optimizer at a 5e-5 learning rate, a batch size of $128$, lora rank of 32, and a $256$ cut-off length.
@@ 4.1 Experimental Results
Vicuna (vicuna, ) $\times$ $\times$ 33.10 40.23 43.69
@@
MathCoder-CL mathcoder $\times$ $\times$ 31.14 11.11 9.22
@@
LLaMa3.1 (llama3, ) $\times$ $\times$ 48.01 49.61 49.32
@@
Table 2: Performance comparison of FineScope-tuned models versus baselines across Pre-algebra (Pre-alg.), Algebra (Alg.), and Counting and Probability (Count.&Prob.) domains.
Model Pruned Tuned Dataset Pre-alg. Alg. Count.&Prob.
Vicuna (vicuna, ) $\times$ $\times$ 14.31 10.17 8.11
@@
MathCoder-CL (mathcoder, ) $\times$ $\times$ 11.60 16.77 13.38
@@
LLaMa3.1 (llama3, ) $\times$ $\times$ 32.77 29.87 20.35
@@
GPT-3 (13B) (gpt3, ) $\times$ $\times$ 6.80 5.30 4.50
GPT-3 (175B) (gpt3, ) $\times$ $\times$ 7.70 6.00 4.70
@@
4.2 Effect of MSDFT and Pruning Dataset
Table 4: Effect of Modified SDFT (MSDFT) and domain-specific pruning methods across domains.
Domain MSDFT Pruning
@@ 4.2 Effect of MSDFT and Pruning Dataset
Modified SDFT Table 4 shows a consistent improvement in performance when modified SDFT is applied, compared to standard fine tuning. Across all three domains, models fine tuned with SDFT outperform their counterparts. For example, we observe performance gains of 1.9% in STEM, 5.8% in Social Sciences, and 10.26% in Humanities. This consistent improvement across diverse domains highlights the effectiveness of a distillation based approach for enhancing model performance in domain specific adaptation. Pruning Dataset Table 4 highlights pruning with the FineScope domain aligned dataset, followed by fine tuning on the same domain data, results in an average accuracy gain of 11.8% compared to pruning with a general purpose corpus (bookcorpus, ). This demonstrates that incorporating domain specific samples during the pruning stage helps retain critical representations that are often lost when using generic data, leading to consistently higher performance across STEM, Social Science, and Humanities benchmarks.

### v3

@@ 3.2.1 Pruning
To enhance model efficiency while retaining domain-specific knowledge, we apply structured pruning using LLM-Pruner (Ma et al., 2023). Rather than relying on general-purpose criteria, we guide the pruning process with a domain-specific dataset $D_{s}$ (as shown in Figure 2c), so that only components relevant and active to the target domain are preserved. Guided by the domain-specific dataset $D_{s}$, the contribution of each model block is estimated using gradient-based attribution. Specifically, we compute an importance score for each block based on the first-order gradient of the task-specific loss with respect to $D_{s}$. This allows the pruning process to preserve components that are most critical for performance in the target domain. (We provide an optimization-based interpretation of domain-conditioned pruning and why conditioning on $D_{s}$ preserves domain-critical components in the supplement; see Section 7.3).
We apply block-wise structured pruning using LLM-Pruner (Ma et al., 2023), which removes redundant components of the model and significantly reduces inference cost. (We compare opther pruning methods under matched pruning ratios in the supplement Section 9.) The original model, denoted by $\mathcal{M}$, is pruned with respect to the domain-specific dataset $D_{s}$ and a pruning ratio $r$, resulting in a compressed model $\mathcal{M}_{r}$:
@@
3.2.2 Fine-Tuning via Teacher-Guided Distillation (TGD)
To mitigate the loss of domain-specific knowledge caused by pruning, we adopt a teacher-guided distillation procedure during fine-tuning. Unlike the original formulation of Self-Distillation Fine-Tuning (SDFT) (Yang et al., 2024), which conditions on intermediate teacher (self-generated) outputs, our approach performs cross-model/unpruned original model knowledge transfer from a teacher model to the pruned student.
In our framework, we adapt SDFT to further refine the pruned model. We generate a distilled dataset using either the original model or a pretrained state-of-the-art teacher, and use it to fine-tune the pruned model. This process helps preserve domain-specific knowledge while improving generalization. In the original SDFT formulation (Yang et al., 2024), given an input $x$, context $c^{t}$, and output $y^{t}$ from the teacher model, the distilled output $y^{\prime}$ is sampled as:
@@ 3.2.2 Fine-Tuning via Teacher-Guided Distillation (TGD)
We adapt this procedure in FineScope by modifying the fine-tuning objective for the pruned model. The resulting distillation loss is defined as:
$L_{\text{tgd}}=-\log f_{p}(y^{\prime}\mid c^{t},y^{t}),$ (16)
where $f_{p}$ denotes the pruned model and $L_{\text{tgd}}$ represents the distillation loss.
This teacher-guided distillation serves two purposes: (1) Knowledge Recovery It transfers domain-specific knowledge that may have been weakened or removed during pruning. (2) Regularization It reduces overfitting to the small curated dataset by exposing the pruned model to high-confidence teacher outputs.
Importantly, the teacher model is used only during distilled dataset construction. The student is trained independently thereafter without direct access to teacher logits or intermediate representations.
@@ 4 Evaluation
(1) Models. We evaluate our method using three models: Vicuna-7B (Zheng et al., 2023), a fine-tuned version of LLaMa 2 (Touvron et al., 2023); MathCoder-CL-7B (Wang et al., 2023), a CodeLlama (Roziere et al., 2023) variant; and LLaMa 3.1-8B (Dubey et al., 2024).
(2) Baselines. We compare FineScope against six baseline settings: (a) fine tuning the pruned model using randomly selected data of the same size, (b) fine tuning with the full dataset containing mixed domains, (c) fine tuning with Alpaca data using FineScope’s pruning strategy, (d) evaluating pretrained models without any fine tuning, (e) evaluating pretrained models fine tuned with FineScope-curated data, and (f) comparisons against GPT-3 (6.7B and 175B)(Brown et al., 2020) and OLMO-7B(Groeneveld et al., 2024) (Table 1), along with GPT-3 (13B and 175B) (Table 2)
@@
(a) Domain Specific Tuning. We prune models using domain specific datasets curated with our SAE-guided framework. The SAEs are trained on the RedPajama dataset (Computer, 2023), which includes content from CommonCrawl, C4, GitHub, Wikipedia, Books3, ArXiv, and StackExchange, providing broad domain coverage. Using these SAEs, we curate domain specific subsets from OpenInstruct (hakurei, 2023), which aggregates instruction datasets such as Alpaca (Lan, 2019), Self Instruct (Wang et al., 2022), GPT 4 Instruct, Roleplay (Teknium, 2023), Code Alpaca (Chaudhary, 2023), and Dolly (Ouyang et al., 2022). Based on user provided seed samples, we extract curated datasets for STEM (2,100 samples), Social Sciences (2,401 samples), and Humanities (2,374 samples), selecting the most frequently chosen samples across all trained SAEs.
(b) Subdomain Specific Tuning (Math). To assess FineScope’s effectiveness at a more granular level, we evaluate fine tuning on mathematical subdomains. SAEs are trained on the MetaMath dataset (Yu et al., 2023), and used to curate subsets from the Math dataset (Hendrycks et al., 2021). In this setting, subdomains are merged into a unified pool for curation. From this pool, we extract the Pre Algebra, Algebra, and Counting and Probability subsets. Models are fine tuned using OpenMath2 (Toshniwal et al., 2024), and Notus 7B (Tunstall et al., 2023) serves as the Alpaca tuned baseline. Evaluations are conducted separately on each subdomain test set (Table 2).
(c) Coding Specific Tuning. Following the same SAE guided curation approach, we construct a code focused dataset from OpenInstruct, resulting in 1,200 examples. To evaluate model performance on code generation, we fine tune on this curated dataset and assess results using the HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) benchmarks (Table 3).
(4) SAE [^1] We used Anthropic style architecture (EleutherAI, 2024). training: We ran SAE training (EleutherAI, 2024) using the AdamW optimizer with a learning rate of 1e-5, a batch size of $8$, and a Top-K value of $128$. We used GPT-4 (Achiam et al., 2023) to generate the $10$ user-defined seeds.
(5) Finetuning: We fine-tuned LMs (LORA-fine tuning (Hu et al., 2021)) using the same AdamW optimizer at a 5e-5 learning rate, a batch size of $128$, lora rank of 32, and a $256$ cut-off length. For generating the distilled dataset, we have used the corresponding unpruned model.
@@ 4.1 Experimental Results
Vicuna (Zheng et al., 2023) $\times$ $\times$ 33.10 40.23 43.69
@@
MathCoder-CL Wang et al. (2023) $\times$ $\times$ 31.14 11.11 9.22
@@
LLaMa3.1 (Dubey et al., 2024) $\times$ $\times$ 48.01 49.61 49.32
@@
Table 2: Performance comparison of FineScope-tuned models versus baselines across Pre-algebra (Pre-alg.), Algebra (Alg.), and Counting and Probability (C.&P.) domains.
Model Pruned Tuned Dataset Pre-alg. Alg. C.&P.
Vicuna (Zheng et al., 2023) $\times$ $\times$ 14.31 10.17 8.11
@@
MathCoder-CL (Wang et al., 2023) $\times$ $\times$ 11.60 16.77 13.38
@@
LLaMa3.1 (Dubey et al., 2024) $\times$ $\times$ 32.77 29.87 20.35
@@
GPT-3 (13B) (Brown et al., 2020) $\times$ $\times$ 6.80 5.30 4.50
GPT-3 (175B) (Brown et al., 2020) $\times$ $\times$ 7.70 6.00 4.70
@@
4.2 Effect of Teacher-Guided Distillation (TGD) and Pruning Dataset
Teacher-Guided Distillation (TGD) Table 4 shows a consistent improvement in performance when modified SDFT is applied, compared to standard fine tuning. Across all three domains, models fine tuned with TGD outperform their counterparts. For example, we observe performance gains of 1.9% in STEM, 5.8% in Social Sciences, and 10.26% in Humanities. This consistent improvement across diverse domains highlights the effectiveness of a distillation based approach for enhancing model performance in domain specific adaptation.
Pruning Dataset Table 4 highlights pruning with the FineScope domain aligned dataset, followed by fine tuning on the same domain data, results in an average accuracy gain of 11.8% compared to pruning with a general purpose corpus (Zhu et al., 2015). This demonstrates that incorporating domain specific samples during the pruning stage helps retain critical representations that are often lost when using generic data, leading to consistently higher performance across STEM, Social Science, and Humanities benchmarks.
Table 4: Effect of Teacher-Guided Distillation(TGD) and domain-specific pruning method across domains.
Domain TGD Pruning
@@ 4.3 Effect of TopK on Computation
4.4 Comparison with Synthetic Dataset
In Figrue 4 we evaluate our curated FineScope dataset (2.1K samples) against the publicly available synthetic STEM-Saraswati dataset (Tiwari, 2024), generated using GPT-4 (Achiam et al., 2023), as well as general-purpose finetuning datasets such as Alpaca and OpenInstruct. scope achieves comparable accuracy to STEM-Saraswati on STEM-specific tasks, demonstrating the high quality of our curated dataset. Moreover, when contrasted with general-purpose datasets, both STEM-Saraswati and FineScope achieve substantially higher performance, highlighting the critical role of high-quality, domain-specific data in enhancing model capabilities within specialized scientific domains.
Figure 4: Performance comparison with synthetic STEM dataset.
4.5 Effect of SAE-Embedding
As shown in Table 5, the method of data curation has a significant impact on model performance after fine-tuning. The results consistently demonstrate that curating the fine-tuning dataset using our interpretable SAE-embeddings leads to substantially better outcomes than using standard raw embeddings. To analyze the effect of the embedding type, we compare the performance of original and pruned models fine-tuned on datasets curated by each method. In addition to BERT-based embeddings, we evaluate a modern sentence embedding model (all-mpnet-base-v2 Sentence Transformers ) and a simple magnitude based top-K selection strategy for SAE training. While sentence transformer embeddings provide moderate improvements over BERT and activation magnitude serves as a lightweight internal baseline, both remain consistently inferior to SAE-Embedding across models and domains. (A broader comparison of Top-K coordinate selectors (random, magnitude, variance, PCA, and ours) is provided in the supplement; see Section 8.)
Using SAE embeddings consistently improves performance across models and domains. For Vicuna, the average gain after fine-tuning is +5.08 for the original model and +2.06 for the pruned model. This suggests that SAE-derived sparse, interpretable features better align with task-relevant representations, leading to higher-quality fine-tuning data.
Table 5: Performance comparison of different dataset selection methods. Sentence-Transformer uses all-mpnet-base-v2. MC-CL: MathCoder-CL.
Model Type Dataset Selection STEM Social Sci. Hum.
Vicuna Full BERT-Embedding 31.17 35.04 34.49
Full Sentence-Transformer 32.08 36.72 36.15
Full SAE-Embedding(Magnitude) 30.92 34.10 33.85
Full SAE-Embedding 33.32 40.21 42.43
Pruned BERT-Embedding 29.01 34.23 34.48
Pruned Sentence-Transformer 29.87 35.02 35.41
Pruned SAE-Embedding(Magnitude) 28.74 33.51 33.02
Pruned SAE-Embedding 31.12 36.23 36.55
MC-CL Full BERT-Embedding 30.45 29.14 29.01
Full Sentence-Transformer 31.82 30.02 29.88
Full SAE-Embedding(Magnitude) 29.73 27.95 27.60
Full SAE-Embedding 34.96 32.91 31.66
Pruned BERT-Embedding 24.91 12.73 12.53
Pruned Sentence-Transformer 25.37 13.11 13.02
Pruned SAE-Embedding(Magnitude) 23.84 11.92 11.74
Pruned SAE-Embedding 25.89 13.81 13.68
LLaMa3.1 Full BERT-Embedding 46.19 49.22 50.94
Full Sentence-Transformer 47.10 50.05 51.21
Full Activation-Magnitude 45.32 47.88 49.03
Full SAE-Embedding 48.84 51.66 51.45
Pruned BERT-Embedding 37.02 39.32 39.00
Pruned Sentence-Transformer 38.11 40.08 40.14
Pruned Activation-Magnitude 36.45 38.02 37.66
Pruned SAE-Embedding 39.91 41.07 41.19